# CooHOI: Learning Cooperative Human-Object Interaction with Manipulated Object Dynamics

### v1

@@ Abstract
Recent years have seen significant advancements in humanoid control, largely due to the availability of large-scale motion capture data and the application of reinforcement learning methodologies. However, many real-world tasks, such as moving large and heavy furniture, require multi-character collaboration. Given the scarcity of data on multi-character collaboration and the efficiency challenges associated with multi-agent learning, these tasks cannot be straightforwardly addressed using training paradigms designed for single-agent scenarios. In this paper, we introduce Cooperative Human-Object Interaction (CooHOI), a novel framework that addresses multi-character objects transporting through a two-phase learning paradigm: individual skill acquisition and subsequent transfer. Initially, a single agent learns to perform tasks using the Adversarial Motion Priors (AMP) framework. Following this, the agent learns to collaborate with others by considering the shared dynamics of the manipulated object during parallel training using Multi-Agent Proximal Policy Optimization (MAPPO). When one agent interacts with the object, resulting in specific object dynamics changes, the other agents learn to respond appropriately, thereby achieving implicit communication and coordination between teammates. Unlike previous approaches that relied on tracking-based methods for multi-character HOI, CooHOI is inherently efficient, does not depend on motion capture data of multi-character interactions, and can be seamlessly extended to include more participants and a wide range of object types.
@@ 1 Introduction
Imagine the scenario where you’re moving home and seeking help from humanoid robots. However, certain items, like beds and sofas, are too large and heavy for a single robot to manage effectively. Despite significant advancements in motion synthesis of physics-based human-object interactions Hassan et al. (2021); Zhang et al. (2022); Pan et al. (2023); Xiao et al. (2023); Merel et al. (2020); Zhao et al. (2023); Hassan et al. (2023) and humanoid robots performing complex operations Seo et al. (2023); Matsuura et al. (2023); Radosavovic et al. (2023); Cheng et al. (2024); Zhang et al. (2024), the area of multiple characters collaboratively transporting objects—particularly when a single individual may find it challenging to handle heavy or bulky items—remains relatively under-investigated. Some previous efforts Zhang et al. (2023) tried to address these challenges using tracking-based methods, capturing interactions between two characters during tasks like box carrying, and training policies to mimic these actions. However, obtaining comprehensive data on interactions involving multiple participants is costly, and tracking-based methods struggle to adapt to scenarios with varying object sizes and an increased number of agents. Another straightforward idea might be to directly train a multi-agent policy from scratch, using an approach similar to that for single agents. However, the expanded action sampling space significantly slows down the overall training process, preventing the policy from converging properly.
@@
Inspired by how humans learn cooperation skills, i.e. beginning with mastering tasks independently and then developing strategies for collaborative efforts in coordination tasks, we propose a two-stage framework for training multi-agent cooperation strategies, named Cooperative Human-Object Interaction, CooHOI. In the initial stage, we train a single-agent object carrying policy utilizing the AMP framework Peng et al. (2021); Hassan et al. (2023) as the foundation for basic policy models. To ensure agents focus significantly on the dynamics of the objects they are tasked with transporting, we include the state and velocity information of the object’s bounding box in the agent’s observations. The second stage aims to transfer these individual carrying skills into collaborative strategies. When two humans collaborate to carry a long object, they typically hold it at opposite ends. Moreover, when one individual takes action, it affects the dynamics of the object, enabling the other individual to perceive this change and adjust their actions accordingly for effective cooperation. Therefore, we adjust the observation for both agents to focus on the bounding boxes located at the ends of the long objects. This refinement aligns with the observation space used during the single-agent training phase, effectively leveraging the previously developed solo carrying skills. Additionally, the inherent rigid body characteristic of long objects facilitates implicit communication between agents. This setup allows an agent to adeptly adjust to their teammate’s actions by observing changes in the object’s dynamics, thereby enhancing coordination and cooperation in carrying tasks. The overall framework is illustrated in Figure 2.
@@ 2 Related Work
Synthesizing natural and physically plausible human-scene interactions, such as humanoid characters sitting on chairs, lying on beds, and carrying boxes, is crucial for advancements in character animation and robotics. Physics-based methods leverage physics simulators Makoviychuk et al. (2021); Todorov et al. (2012) to control characters modeled as interconnected rigid bodies through joint torques and deep reinforcement learning methods. To facilitate the training process, some strategies employ tracking-based techniques Liu et al. (2015); Peng et al. (2018); Chao et al. (2021); Xie et al. (2023); Merel et al. (2020); Zhang et al. (2023), which rely on the availability of high-quality reference motions. This reliance often limits their application in human-scene interactions due to the scarcity of suitable data and affects their versatility across different scenarios. Recently, the Adversarial Motion Priors(AMP) framework Peng et al. (2021) introduced the use of a discriminator to ensure that generated motions align with the distribution of reference motions. This approach has shown success in various downstream tasks Juravsky et al. (2022); Peng et al. (2022); Tessler et al. (2023), including human-scene interactions Hassan et al. (2023); Pan et al. (2023); Xiao et al. (2023). Despite these advancements, there has been limited focus on synthesizing cooperative behaviors among multiple characters interacting with objects—a gap that our work aims to address.
@@
While significant advancements have been made in synthesizing motions for single agents, the realm of multi-character animation remains relatively unexplored. Existing approaches predominantly rely on kinematic-based methods Won et al. (2014); Liang et al. (2023); Tanaka and Fujiwara (2023); Shafir et al. (2023); Ghosh et al. (2023) and datasets of multi-character interactions Kundu et al. (2020); Shen et al. (2019); Fieraru et al. (2020). These methods, however, require high-quality interaction data and often fall short in ensuring physical plausibility or adequately handling multi-character cooperative interactions with objects. Physics-based techniques for character animation have primarily focused on aspects like crowd navigation Haworth et al. (2020); Rempe et al. (2023), limiting themselves to behaviors such as pedestrian movement and collision avoidance. Although  Zhang et al. (2023) showcases interactions among multiple characters and object manipulation, such as carrying long boxes, these tracking-based approaches necessitate motion capture data of human interactions or human-object interactions and struggle to scale to an arbitrary number of agents or different types of objects. In contrast, our framework requires only single agent motion capture data for multi-character object transporting tasks and can easily extend to different types of objects and different numbers of agents.
Figure 2: Our framework employs a two-phase learning paradigm. In the first phase, depicted on the left, we train single-agent carrying skills using the AMP framework. In the second phase, we transfer these single-agent skills to a cooperative context. Notably, we use the dynamics of the object as feedback information, as illustrated by the bounding box shown in the figures.
@@ 3.1 Preliminary
We formulate physics-based character control as a goal-conditioned reinforcement learning task. At each time step $t$, the agent samples an action from its policy $\pi\left(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t}\right)$ based on the current state $\mathbf{s}_{t}$ and the task-specific goal feature $\mathbf{g}_{t}$. When this action is applied to the character, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives a task reward $r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)$. To train control policies that enable characters to achieve high-level tasks in a natural and life-like manner, we adopt the AMP framework Peng et al. (2021). While this framework aims to optimize the expected cumulative task reward $J(\pi)$, it introduces a discriminator to encourage the character to produce behaviors similar to those in the dataset by providing a style reward $r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right)$. The agent’s reward $r_{t}$ at each time step $t$ is defined by $r_{t}=w^{G}r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)+w^{S}r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right)$. More details can be found in our appendix and the original AMP paper Peng et al. (2021).
@@
We formulate our cooperative task as a Partially Observable Markov Decision Process (POMDP) Littman (1994). A POMDP with $n$ agents is defined by $\left\{\mathcal{S},\mathcal{A}_{1},\cdots,\mathcal{A}_{n},\mathcal{O}_{1},\cdots,\mathcal{O}_{n},\mathcal{R},\mathcal{P},\mathcal{T}\right\}$, where $\mathcal{S}$ is the state space, $\mathcal{A}_{i}$ is the action space for agent $i$, $O_{i}$ is the local observation for agent $i$, and $\mathcal{R}$ is the shared reward function. Each agent uses its own policy $\pi_{\theta}\left(\mathbf{a}_{i}\mid\mathbf{o}_{i}\right)$ to take action $\mathbf{a}_{i}\in\mathcal{A}_{i}$ based on its local observation $\mathbf{o}_{i}\in O_{i}$. The environment transitions according to the function $\mathcal{P}\left(\mathbf{s}_{t+1}\mid\mathbf{s}_{t},\mathbf{a}_{1},\cdots,\mathbf{a}_{n}\right)$, where $\mathbf{s}_{t}$, $\mathbf{s}_{t+1}$ are states of time step $t$ and $t+1$, respectively. The agents then get a reward $\mathbf{r}_{t}$ based on the states $\mathbf{s}_{t}$ and $\mathbf{s}_{t+1}$. The goal of multi-agent reinforcement learning algorithms is to jointly optimize the discounted accumulated reward $J(\theta)=\mathbb{E}_{\mathbf{a}_{1}^{t},\cdots,\mathbf{a}_{n}^{t},s^{t}}\left[\Sigma_{t=0}^{T}\gamma^{t}\mathbf{r}_{t}\right]$.
@@ 3.2 Single Agent Carrying Skills Training
In developing our approach for single-agent object manipulation, we introduce several advancements based on previous methods. We integrate the dynamics of the manipulated object into the observation and introduce a reward function framework for object manipulation tasks. These enhancements allow the trained policy to easily adapt to multi-agent settings.
@@
By incorporating these dynamics into the observation, we establish a feedback mechanism that keeps agents continually informed about the outcomes of their actions. This also equips agents with the capability to react appropriately, whether engaged in single-agent tasks or collaborative multi-agent environments. Along with the state of the agent $s_{t}$ and the position of the target $d_{t}^{\mathrm{pos}}$, we formulate the dynamics-aware task observation as:
@@ 3.3 Cooperation Strategy Training
Upon mastering single-agent tasks involving object carrying, it is crucial to effectively transfer and further enhance these abilities to boost the cooperative learning process. Initially, we facilitate the skill transfer by guiding each agent to lift and transport one end of a large object. Subsequently, we refine the control policy through the application of the Multi-Agent Proximal Policy Optimization (MAPPO) algorithm Yu et al. (2022), leveraging the dynamics of manipulated objects for feedback and implicit communication.
@@
Moreover, this design of dynamics as observation acts as implicit communication between agents. When an agent takes an action, the object dynamics will change, and because we use the manipulated object dynamics as feedback in single-agent training, the change in object dynamics will result in a corresponding change in strategy for the other agents. This method of implicit communication presents a straightforward yet effective approach for enhancing teamwork in multi-agent settings. Furthermore, this framework proves to be adaptable with variations in the number of participating agents.
3.3.2 Cooperation Training using MAPPO Algorithm.
The non-stationary environment of multi-agent reinforcement learning, coupled with sample efficiency issues, presents a significant challenge in training agents to collaborate effectively. During the cooperation training phase, we continue to use a reward function similar to that used in the single-agent training phase and employ the centralized training and decentralized execution (CTDE) Littman (1994) scheme for coordination training. During the training phase, we utilize the Multi-Agent Proximal Policy Optimization (MAPPO) algorithm Yu et al. (2022) to develop cooperative strategies among agents. This approach involves updating the value function network $V_{\phi}$ according to Equation (6), using the trajectories $\mathcal{D}$ that are accumulated and shared among all agents. More details can be found in our appendix and the original MAPPO paper Yu et al. (2022).
@@
In addition, given the homogeneous roles of agents in the collaborative task of carrying long objects, we adopt the strategy of parameter sharing, a method proven to enhance performance across various cooperative tasks Christianos et al. (2021); Terry et al. (2020); Yu et al. (2022). Specifically, we share the parameters of both the policy and value networks among all agents to improve the training of cooperative behaviors.
@@ 4 Experiments
We conducted extensive experiments to test the effectiveness and also the boundary of capabilities of our framework. The basic experiment setups are explained in Section 4.1 and in Appendix. We evaluate our framework on various object-carrying tasks in Section 4.2. To better understand the importance of different design decisions in our framework, we performed extensive ablation studies in Section 4.3. Since our method primarily focuses on interactions between characters and objects, we also provide extensive visual analysis and presentations to demonstrate our framework. Please refer to the supplementary video for more details.
@@ 4.1 Experiments Setup
Our primary source of motion data is the AMASS dataset Mahmood et al. (2019), which provides motions encoded in SMPL Pavlakos et al. (2019) parameters. We collected a total of four types of basic reference motion data, including 9 motions related to walking, 5 related to picking up, 4 related to carrying, and 5 related to putting down. To enhance the robustness of the carrying process, we randomly initialized the weight and size of the object, as well as the distance from the person to the destination. Specifically, for a single individual, the object’s weight ranged from 5KG to 25KG, its size varied between 0.5 to 1.5 times its original scale, and the distances between the agent and the object, as well as between the object and the destination, ranged from 1 m to 20 m. To enhance the robustness of the carrying process, we randomly varied the weight, size, and of the object, as well as the distance from the person to the endpoint. Furthermore, considering that the process of multiple people carrying might involve situations where someone walks backward, we ensured that the single agent mastered the basic movements for multi-person collaboration by introducing additional motion data involving backward and side walking. More training details can be found in the appendix.
Table 1: This table presents our results for single-agent and two-agent carrying of the Box object. “Ours” refers to the policy trained using the complete CooHOI framework in both single-agent and two-agent settings. “Ours+WeightAug” indicates that we applied the same weight augmentation design as InterPhys Hassan et al. (2023).
@@
Single Agent InterPhys Hassan et al. (2023) [5,26] [1,10] 94.3 8.3
Ours+WeightAug [2,26] [1,20] 93.98 4.8
Ours [2,13] [1,20] 96.48 6.9
Two Agent w/o CooHOI [15,40] [2,20] 0 NAN
Ours [15,40] [2,20] 89.54 3.86
@@
Following InterPhys Hassan et al. (2023), we utilize Success Rate and Precision as the primary metrics. Specifically, a task is considered successful if the distance between the object and its destination is less than 20cm. Precision measures the distance between the object and the target point across all completed tasks. Moreover, all evaluations are based on average results across 4096 environments and 4 random seeds.

### v2

@@ Abstract
Enabling humanoid robots to clean rooms has long been a pursued dream within humanoid research communities. However, many real-world tasks, such as moving large and heavy furniture, require multi-humanoid collaboration. Given the scarcity of motion capture data on multi-humanoid collaboration and the efficiency challenges associated with multi-agent learning and control, these tasks cannot be straightforwardly addressed using training paradigms designed for single-agent scenarios. In this paper, we introduce Cooperative Human-Object Interaction (CooHOI), a framework designed to tackle the challenge of multi-humanoid object transportation problem through a two-phase learning paradigm: individual skill learning and subsequent policy transfer. First, a single humanoid character learns to interact with objects through imitation learning from human motion priors. Then, the humanoid learns to collaborate with others by considering the shared dynamics of the manipulated object using centralized training and decentralized execution (CTDE) multi-agent RL algorithms. When one agent interacts with the object, resulting in specific object dynamics changes, the other agents learn to respond appropriately, thereby achieving implicit communication and coordination between teammates. Unlike previous approaches that relied on tracking-based methods for multi-humanoid HOI, CooHOI is inherently efficient, does not depend on motion capture data of multi-humanoid interactions, and can be seamlessly extended to include more participants and a wide range of object types.
@@ 1 Introduction
Imagine the scenario where you’re moving home and seeking help from humanoid robots. However, certain items, like beds and sofas, are too large and heavy for a single robot to manage effectively. Despite significant advancements in learning and control of physics-based human-object interactions [6, 43, 21, 37, 20, 45, 7] and humanoid robots performing complex operations [29, 19, 26, 2, 42, 41, 10, 9, 39], the area of multiple humanoids collaboratively transporting objects—particularly when a single individual may find it challenging to handle heavy or long items—remains relatively under-investigated. Some previous efforts [44] tried to address these challenges using tracking-based methods, capturing interactions between two characters during tasks like box carrying, and training policies to mimic these actions. However, obtaining comprehensive motion data on interactions involving multiple participants is costly, and tracking-based methods struggle to adapt to scenarios with varying object sizes and an increased number of agents. Another straightforward idea might be to directly train a multi-agent policy from scratch, using an approach similar to that for single agents. However, the expanded action sampling space significantly slows down the overall training process, preventing the policy from converging properly.
@@
Inspired by how humans learn cooperation skills, i.e. beginning with mastering tasks independently and then developing strategies for collaborative efforts in coordination tasks, we propose a two-stage framework for training multi-agent cooperation strategies, named Cooperative Human-Object Interaction, CooHOI. In the initial stage, we train a single-agent object carrying policy utilizing the AMP framework [25, 7]. To ensure agents focus significantly on the dynamics of the objects they are tasked with transporting, we include the state and velocity information of the object’s bounding box in the agent’s observation space. The second stage aims to transfer these individual carrying skills into collaborative strategies. When two humans collaborate to carry a long object, they typically hold it at opposite ends. Moreover, when one individual takes action, it affects the dynamics of the object, enabling the other individual to perceive this change and adjust their actions accordingly for effective cooperation. Therefore, we adjust the observation for both agents to focus on the bounding boxes located at the ends of the long objects. This refinement aligns with the observation space used during the single-agent training phase, effectively leveraging the previously developed single carrying skills. Additionally, the inherent rigid body characteristic of long objects facilitates implicit communication between agents. This setup allows an agent to adeptly adjust to their teammate’s actions by observing changes in the object’s dynamics, thereby enhancing coordination and cooperation in carrying tasks. The overall framework is illustrated in Figure 2.
@@ 2 Related Work
Synthesizing natural and physically plausible human-scene interactions, such as humanoid characters sitting on chairs, lying on beds, and carrying boxes, is crucial for advancements in character animation and robotics. Physics-based methods leverage physics simulators [18, 35] to control characters modeled as interconnected rigid bodies through joint torques and deep reinforcement learning methods. To facilitate the training process, some strategies employ tracking-based techniques [16, 23, 1, 38, 20, 44], which rely on the availability of high-quality reference motions. This reliance often limits their application in human-scene interactions due to the scarcity of suitable data and affects their versatility across different scenarios. Recently, the Adversarial Motion Priors(AMP) framework [25] introduced the use of a discriminator to ensure that generated motions align with the distribution of reference motions. This approach has shown success in various downstream tasks [11, 24, 34], including human-scene interactions [7, 21, 37]. Despite these advancements, there has been limited focus on synthesizing cooperative behaviors among multiple characters interacting with objects—a gap that our work aims to address.
@@
While significant advancements have been made in synthesizing motions for single agents, the realm of multi-character animation remains relatively unexplored. Existing approaches predominantly rely on kinematic-based methods [36, 14, 32, 30, 5] and datasets of multi-character interactions [13, 31, 4]. These methods, however, require high-quality interaction data and often fall short in ensuring physical plausibility or adequately handling multi-character cooperative interactions with objects. Physics-based techniques for character animation have primarily focused on aspects like crowd navigation [8, 27], limiting themselves to behaviors such as pedestrian movement and collision avoidance. Although  [44] showcases interactions among multiple characters and object manipulation, such as carrying long boxes, these tracking-based approaches necessitate motion capture data of human interactions or human-object interactions and struggle to scale to an arbitrary number of agents or different types of objects. In contrast, our framework requires only single agent motion capture data for multi-character object transporting tasks and can easily extend to different types of objects and different numbers of agents.
Figure 2: Our framework employs a two-phase learning paradigm. In the first phase, depicted on the left, we train single-agent carrying skills by imitating from human motion priors. In the second phase, we transfer these single-agent skills to a cooperative context. Notably, we use the dynamics of the object as feedback information, as illustrated by the bounding box shown in the figures.
@@ 3.1 Preliminary
We formulate physics-based character control as a goal-conditioned reinforcement learning task. At each time step $t$, the agent samples an action from its policy $\pi\left(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t}\right)$ based on the current state $\mathbf{s}_{t}$ and the task-specific goal feature $\mathbf{g}_{t}$. When this action is applied to the character, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives a task reward $r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)$. To train control policies that enable characters to achieve high-level tasks in a natural and life-like manner, we adopt the AMP framework [25]. While this framework aims to optimize the expected cumulative task reward $J(\pi)$, it introduces a discriminator to encourage the character to produce behaviors similar to those in the dataset by providing a style reward $r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right)$. The agent’s reward $r_{t}$ at each time step $t$ is defined by $r_{t}=w^{G}r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)+w^{S}r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right)$. More details can be found in our appendix and the original AMP paper [25].
@@
We formulate our cooperative task as a Partially Observable Markov Decision Process (POMDP) [15]. A POMDP with $n$ agents is defined by $\left\{\mathcal{S},\mathcal{A}_{1},\cdots,\mathcal{A}_{n},\mathcal{O}_{1},\cdots,\mathcal{O}_{n},\mathcal{R},\mathcal{P},\mathcal{T}\right\}$, where $\mathcal{S}$ is the state space, $\mathcal{A}_{i}$ is the action space for agent $i$, $O_{i}$ is the local observation for agent $i$, and $\mathcal{R}$ is the shared reward function. Each agent uses its own policy $\pi_{\theta}\left(\mathbf{a}_{i}\mid\mathbf{o}_{i}\right)$ to take action $\mathbf{a}_{i}\in\mathcal{A}_{i}$ based on its local observation $\mathbf{o}_{i}\in O_{i}$. The environment transitions according to the function $\mathcal{P}\left(\mathbf{s}_{t+1}\mid\mathbf{s}_{t},\mathbf{a}_{1},\cdots,\mathbf{a}_{n}\right)$, where $\mathbf{s}_{t}$, $\mathbf{s}_{t+1}$ are states of time step $t$ and $t+1$, respectively. The agents then get a reward $\mathbf{r}_{t}$ based on the states $\mathbf{s}_{t}$ and $\mathbf{s}_{t+1}$. The goal of multi-agent reinforcement learning algorithms is to jointly optimize the discounted accumulated reward $J(\theta)=\mathbb{E}_{\mathbf{a}_{1}^{t},\cdots,\mathbf{a}_{n}^{t},s^{t}}\left[\Sigma_{t=0}^{T}\gamma^{t}\mathbf{r}_{t}\right]$.
@@ 3.2 Single Agent Carrying Skills Training
In developing our approach for single-agent object manipulation, we introduce several advancements based on previous methods. We integrate the dynamics of the manipulated object into the observation space and introduce a reward function framework for object manipulation tasks. These enhancements allow the trained policy to easily adapt to multi-agent settings.
@@
By incorporating these dynamics information into the observation, we establish a feedback mechanism that keeps agents continually informed about the outcomes of their actions. This also equips agents with the capability to react appropriately, whether engaged in single-agent tasks or collaborative multi-agent environments. Along with the state of the agent $s_{t}$ and the position of the target $d_{t}^{\mathrm{pos}}$, we formulate the dynamics-aware task observation as:
@@ 3.3 Cooperation Strategy Training
Upon mastering single-agent tasks involving object carrying, it is crucial to effectively transfer and further enhance these abilities to boost the cooperative learning process. Initially, we facilitate the skill transfer by guiding each agent to lift and transport one end of a large object. Subsequently, we refine the control policy through the application of the Multi-Agent Proximal Policy Optimization (MAPPO) algorithm [40], leveraging the dynamics of manipulated objects for feedback and implicit communication.
@@
Moreover, this design of dynamics as observation acts as implicit communication channel between agents. When an agent takes an action, the object dynamics will change, and because we use the manipulated object dynamics as feedback in single-agent training, the change in object dynamics will result in a corresponding change in strategy for the other agents. This method of implicit communication presents a straightforward yet effective approach for enhancing teamwork in multi-agent settings. Furthermore, this framework proves to be adaptable with variations in the number of participating agents.
3.3.2 Cooperation Training using CTDE Scheme.
The non-stationary environment of multi-agent reinforcement learning, coupled with sample efficiency issues, presents a significant challenge in training agents to collaborate effectively. During the cooperation training phase, we continue to use a reward function similar to that used in the single-agent training phase and employ the centralized training and decentralized execution (CTDE) [15] scheme for coordination training. During the training phase, we utilize the Multi-Agent Proximal Policy Optimization (MAPPO) algorithm [40] to develop cooperative strategies among agents. This approach involves updating the value function network $V_{\phi}$ according to Equation (6), using the trajectories $\mathcal{D}$ that are accumulated and shared among all agents. More details can be found in our appendix and the original MAPPO paper [40].
@@
In addition, given the homogeneous roles of agents in the collaborative task of carrying long objects, we adopt the strategy of parameter sharing, a method proven to enhance performance across various cooperative tasks [3, 33, 40]. Specifically, we share the parameters of both the policy and value networks among all agents to improve the training of cooperative behaviors.
@@ 4 Experiments
We conducted extensive experiments to test the effectiveness and also the boundary of capabilities of our framework. The basic experiment setups are explained in Section 4.1 and in Appendix. We evaluate our framework on various object-carrying tasks in Section 4.2. To better understand the importance of different design decisions in our framework, we performed extensive ablation studies in Section 4.3. Since our method primarily focuses on interactions between characters and objects, we also provide extensive visual analysis and presentations to demonstrate our framework.
@@ 4.1 Experiments Setup
Our primary source of motion data is the AMASS dataset [17], which provides motions encoded in SMPL [22] parameters. We collected a total of four types of basic reference motion data, including 9 motions related to walking, 5 related to picking up, 4 related to carrying, and 5 related to putting down. To enhance the robustness of the carrying process, we randomly initialized the weight and size of the object, as well as the distance from the person to the destination. Specifically, for a single individual, the object’s weight ranged from 5KG to 25KG, its size varied between 0.5 to 1.5 times its original scale, and the distances between the agent and the object, as well as between the object and the destination, ranged from 1 m to 20 m. To enhance the robustness of the carrying process, we randomly varied the weight, size, and of the object, as well as the distance from the person to the endpoint. Furthermore, considering that the process of multiple people carrying might involve situations where someone walks backward, we ensured that the single agent mastered the basic movements for multi-person collaboration by introducing additional motion data involving backward and side walking. More training details can be found in the appendix.
Table 1: This table presents our results for single-agent and two-agent carrying of the Box object. “CooHOI” refers to the policy trained using the complete CooHOI framework in both single-agent and two-agent settings. “CooHOI+WeightAug” indicates that we applied the same weight augmentation design as InterPhys [7].
@@
Single Agent InterPhys [7] [5,26] [1,10] 94.3 8.3
CooHOI+WeightAug(Ours) [2,26] [1,20] 93.98 4.8
CooHOI(Ours) [2,13] [1,20] 96.48 6.9
Two Agent From Scratch [15,40] [2,20] 0 NAN
CooHOI(Ours) [15,40] [2,20] 89.54 3.86
@@
Following InterPhys [7], we utilize Success Rate and Precision as the primary metrics. Specifically, a task is considered successful if the distance between the object and its destination is less than 0.2m. Precision measures the distance between the object and the target point across all completed tasks. Moreover, all evaluations are based on average results across 4096 environments and 4 random seeds.